# Data Mixing Agent: Learning to Re-weight Domains for Continual Pre-training

### v1

@@ 5.1 Continual Pre-training
5.2 Data Re-weighting in Pre-training
@@
Recent works have also explored the effect of domain space definition on data reweighting performance (Wettig et al., 2025; Rukhovich et al., 2025; Diao et al., 2025; Xi et al., 2025), strengthening the importance of carefully defined domains. For example, previous data mixing methods mostly utilized the default domain space defined by data sources. Wettig et al. (2025) carefully defined a 24-dimensional domain space from both the topic (e.g., Science&\&&Tech, Fashion&\&&Beauty) and format (e.g., Academic writing, Content listing) perspectives, and re-organized the training data into these domain spaces. Extensive data mixing experiments on these novel domain spaces showed their effectiveness in improving model training performances compared to the source-based domain space. Inspired by their success, we also train the Data Mixing Agent based on these superior ways of domain space definition.
6 Conclusion
In this paper, we propose the Data Mixing Agent, the first model-based domain reweighting method for continual pre-training, which learns general heuristics for balancing model capabilities on multiple target fields via randomly sampled data mixing trajectories and feedback from an evaluation environment. Extensive experiments show that the agent significantly outperforms strong baseline methods in overall results on 12 general and math reasoning benchmarks. The learned heuristics also generalize well across source-field data, target models, domain spaces, and new target fields such as code generation, without retraining the agent model. Further analysis showcases the data mixing agents’ well-aligned heuristics with human intuitions and their efficiency in achieving superior performance in the target fields with less source-field data.

### v2

@@ F.1 Continual Pre-training
F.2 Data Re-weighting in Pre-training
@@
Recent works have also explored the effect of domain space definition on data reweighting performance (Wettig et al., 2025; Rukhovich et al., 2025; Diao et al., 2025; Xi et al., 2025), strengthening the importance of carefully defined domains. For example, previous data mixing methods mostly utilized the default domain space defined by data sources. Wettig et al. (2025) carefully defined a 24-dimensional domain space from both the topic (e.g., Science&\&Tech, Fashion&\&Beauty) and format (e.g., Academic writing, Content listing) perspectives, and re-organized the training data into these domain spaces. Extensive data mixing experiments on these novel domain spaces showed their effectiveness in improving model training performances compared to the source-based domain space. Inspired by their success, we also train the Data Mixing Agent based on these superior ways of domain space definition.